OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that contains the points (4, -7) and (0, 5).

OpenStudy (anonymous):

y - (-7) = -7 -5 (x - 4 ) ------ 4 - 0 y + 7 = -3x + 12 y = -3x +5